Consider a market where supply and demand are given by Qx s = -16+Px and Qd=92-2Px suppose the gouvernment imposes a price floor of $40, and agrees to purchase any and all units consumers do not buy at the floor price of $40 per unit. determine the cost to the gouvernment of buying firm's unsold units. Compoute the lost social welfare (deadweight loss) that stems from the $40 price floor


1
Expert's answer
2022-02-16T13:13:57-0500

Government cost of unsold units

"Q_S=P_x-16"

"Q_d=92-2P_X"

Initial equilibrium points:

"P_x-16=92-2P_x"

"P_x+2P_x=92+16"

"3P_x=108"

"P_x=36" "Q_S\to 36-16=20" "Q_d\to 92-2(36)=20"

Price Floor "(P_f)=\$ 480"

"Q_s=40-16=24"

"Q_d=92-2(40)=12"

Unsold units = "24-12=12"

Cost of unsold units= "12 \times 40=\$ 480"

Cost of unsold units= Government cost = $ 480


Dead-weight loss

="(Q_d\ at P_x=36)- (Q_d\ at\ P_x =40)"

= "20-10=10"

cost = "10\times 40=\bold{\$400}"